OpenStudy (anonymous):

Use the Pythagorean Theorem to determine which of the following give the measures of the legs and hypotenuse of a right triangle. Check all that apply. 3, 4, 5 4, 11, 14 9, 14, 17 8, 14, 16 8, 15, 17

OpenStudy (dumbcow):

a^2 +b^2 = c^2, where c is the hypotenuse or longest side plug in the numbers and see if they fit in the equation

OpenStudy (anonymous):

For such sums find the square of all the three numbers and then see if the sum of the squares of the smaller two numbers is equal to the square of the largest number or not. If it is, then they form a Pythagorean triplet otherwise not Take the first case 3^2 = 9 4^2 = 16 5^2 = 25 We find 9 + 16 = 25 i.e. 3^2 + 4^2 = 5^2 so they form a Pythagorean triplet and hence can be the legs and hypotenuse of a right angles triangle

OpenStudy (anonymous):

You can follow the same procedure for other sets of numbers

OpenStudy (anonymous):

thank you got the answer